# Very Massive Stars and High N/O: A Tale of the Nitrogen-enriched Super Star Cluster in the Sunburst Arc

### v1

@@ 6 Conclusion
The process studied here may be related to the problem of multiple stellar populations in globular clusters, i.e., significant star-to-star variations in abundances of light elements like He, C, N, O, Na, and Al (Bastian & Lardo, 2018; Milone & Marino, 2022). We note that stellar winds from VMSs, as described above, can mix effectively with the pristine gas in the deep potential well of the super star cluster, and the mixture can give birth to a second generation of stars, leaving a unique chemical imprint.
There are several caveats for this study. For example, the PARSEC data used here are based on non-rotating stars (Costa et al., 2025), while stellar rotation and other stellar physics details may lead to uncertainties in the sub-grid model that we employed. The study also adopted FIRE-3 physics with a presumed IMF for a trade-off in computational costs, while more self-consistent simulations like STARFORGE (Grudić et al., 2021) are desired for a more detailed study of super star cluster formation. For example, the spectroscopically implied number of VMSs falls short of the prediction from the standard IMF (Meštrić et al., 2023), possibly suggesting the importance of a more realistic IMF in these environments, which could be addressed with high-resolution simulations (it may also be a result of the rapid evolution of VMSs). Another shortcoming of our simulations that would ideally require an even higher spatial resolution to improve is the interaction between the hot shocked wind gas and the cool photo-ionized gas, through turbulent mixing and cooling (e.g., Ji et al., 2019), which may be important for the dynamics of the enriched dense gas near the cluster center. There are additional physics that may be important in newborn super star clusters, such as runaway stellar mergers, which we do not model. We leave these investigations for future work.
@@
Y.S., N.M., and C.S.Y. acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) under the funding reference number 568580. L.D. acknowledges research grant support from the Alfred P. Sloan Foundation (Award Number FG-2021-16495) and from the Frank and Karen Dabby STEM Fund in the Society of Hellman Fellows. Part of the computation is performed at the University of Toronto cluster “Niagara,” supported by SciNet (scinethpc.ca) and the Digital Research Alliance of Canada (alliancecan.ca).

### v2

@@ VI Conclusion
The process studied here may be related to the problem of multiple stellar populations in globular clusters, i.e., significant star-to-star variations in abundances of light elements like He, C, N, O, Na, and Al (BastianLardo_2018ARA&A..56...83B; MiloneMarino_2022Univ....8..359M). We note that stellar winds from VMSs, as described above, can mix effectively with the pristine gas in the deep potential well of the super star cluster, and the mixture can give birth to a second generation of stars, leaving a unique chemical imprint.
There are several caveats to this study. The PARSEC tracks used here assume non-rotating stars (CostaShepherdBressan_2025A&A...694A.193C), while stellar rotation introduces uncertainties in the yield model. textcolorblackThe detailed physics of VMS modeling, like the mass loss recipes, can be complicated and metallicity-dependent (SanderVink_2020MNRAS.499..873S; SabhahitVinkSander_2023MNRAS.524.1529S). Another uncertainty is the VMS wind velocity, since observations found a fast terminal velocity of $\sim 3000\,\rm km/s$ for VMSs in the R136 clusters (BrandsdeKoterBestenlehner_2022A&A...663A..36B), unlike the slow-wind model (Vink_2018A&A...615A.119V) used here. All these factors may change the result of the present study. The simulations employ FIRE-3 physics with a presumed IMF for computational efficiency, whereas fully self-consistent frameworks such as STARFORGE (GrudicGuszejnovHopkins_2021MNRAS.506.2199G) would better capture the formation of single stars. Observations suggest that the number of VMSs inferred spectroscopically is lower than predicted by a standard IMF (MestricVanzellaUpadhyaya_2023A&A...673A..50M), highlighting the need for a more realistic IMF in these environments (or reflecting the rapid evolution of VMSs). Our simulations also lack the resolution to model turbulent mixing and cooling between hot shocked winds and cool photoionized gas (e.g., JiOhMasterson_2019MNRAS.487..737J), which may affect the dynamics of enriched dense gas near cluster centers. Additional physics, such as runaway stellar mergers, are likewise omitted. We leave these to future work.
@@
We thank Kendall Shepherd and Emil Rivera-Thorsen for useful comments. Y.S., N.M., and C.S.Y. acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) under the funding reference number 568580. L.D. acknowledges research grant support from the Alfred P. Sloan Foundation (Award Number FG-2021-16495) and from the Frank and Karen Dabby STEM Fund in the Society of Hellman Fellows. Part of the computation is performed at the University of Toronto cluster “Niagara,” supported by SciNet (scinethpc.ca) and the Digital Research Alliance of Canada (alliancecan.ca).